Find the linear velocities and accelerations of centers of sphere, disc and hoop that roll down an inclined plane without slipping. The incline of height h =1m makes an angle of 300 to the horizontal. The initial velocity of all objects v0 =0, compare calculated velocities and accelerations with the velocity and acceleration of the box, which slides from this incline without friction.


1
Expert's answer
2021-03-23T11:15:55-0400

Explanations & Calculations

  • By applying F=ma along the moving direction & using "\small \tau=I\alpha" simultaneously (as the rolling object does not slip, friction helps it roll by the generated torque), linear acceleration can be found & then using "\small V^2=U^2+2as", linear velocity can be found as the linear acceleration is already found & it is constant for the given object.
  • The length of the incline is

"\qquad\qquad
\begin{aligned}
\small s&=\small \frac{1}{\sin30}=\bold{2m}
\end{aligned}"

  • For the sphere (a hollow sphere)

"\qquad\qquad
\begin{aligned}
\small I_s&=\small \frac{2mr^2}{3}\\
\small mg\sin\theta-f&=\small ma\cdots(1)\\
\small f\cdot r&=\small \frac{2mr^2}{3}\cdot\frac{a}{r}\\
\small f&=\small \frac{2ma}{3}\cdots(2) \\
\small\text{By (1) and (2)}\\
\therefore \small a_s&=\small \frac{3}{5}\cdot g\sin\theta=\bold{2.94ms^{-2}}\\
\\
\small V_s&=\small \sqrt{\frac{6g}{5}}=\bold{3.429ms^{-1}}
\end{aligned}"

  • For the disc

"\qquad\qquad
\begin{aligned}
\small I_d&=\small \frac{mr^2}{2}\\
\small mg\sin\theta-f&=\small ma\cdots(1)\\
\small f\cdot r&=\small \frac{mr^2}{2}\cdot\frac{a}{r}\\
\small f&=\small \frac{ma}{2}\cdots(2) \\
\small\text{By (1) and (2)}\\
\therefore \small a_d&=\small \frac{2}{3}\cdot g\sin\theta=\bold{3.26ms^{-2}}\\
\\
\small V_d &=\small \sqrt{\frac{4g}{3}}=\bold{3.615ms^{-1}}
\end{aligned}"

  • For the hoop

"\qquad\qquad
\begin{aligned}
\small I_h&=\small mr^2\\
\\
\small mg\sin\theta-f&=\small ma\cdots(1)\\
\small f\cdot r&=\small mr^2\cdot\frac{a}{r}\\
\small f&=\small ma\cdots(2) \\
\small\text{By (1) and (2)}\\
\therefore \small a_h&=\small \frac{1}{2}\cdot g\sin\theta=\bold{2.45ms^{-2}}\\
\\
\small \text{Then,}\\
\small V_h&=\small \sqrt{g}=\bold{3.13ms^{-1}}
\end{aligned}"

  • For the block

"\qquad\qquad
\begin{aligned}
\small mg\sin\theta&=\small ma\\
\small a_b&=\small g\sin\theta=\bold{4.9ms^{-2}}\\
\\
\small \text{then,}
\\
\small V_b&=\small \sqrt{4g\sin\theta}=\bold{4.43ms^{-1}}
\end{aligned}"


  • Comparison

"\qquad\qquad
\begin{aligned}
\small I_b<I_d&<I_s<I_h\\
\small a_b>a_d&>a_s>a_h\\
\small V_b>V_d&>V_s>V_h
\end{aligned}"